iosangel:

x^2+11x+28=0 how to solve by factoring?

mhchen:

Can you think of 2 numbers that multiply to 28, and add up to 11?

RyeBread1:

i know what multiplies to 28, that is 7x4=28 but do those same numbers have to add ,7 and 4, up to 11 or just any numbers that add up to 11

mhchen:

7 and 4 nice x^2 + 11x + 28 = 0 x^2 + 7x + 4x + 28 = 0 x(x+7)+4(x+7)=0 (x+4)(x+7) = 0

RyeBread1:

so whats the next step after that

iosangel:

Then I would write that as an answer? @mhchen

mhchen:

uh.... you know anything times 0 is 0. so see what happens when x is -4 or -7